In three years, Chad will be three times my present age. I will then be half as old as he. How old am I?
I think it should be set up like this: 3x+3=1/2c
"x" for present age, "c" for chad's age
i know this can't be right because i can't get an answer that makes since.

Answer by fractalier(6550)

When you do age problems, always set up the variables NOW...
Let x be your age now and c be Chad's age now. Thus
c + 3 = 3x and
x + 3 = (1/2)(c + 3)
now double the second equation and substitute and we get
2(x + 3) = c + 3 so that
2(x + 3) = 3x
2x + 6 = 3x
x = 6
c = 15